OpenStudy (anonymous):

Exactly how many cubic inches of ice cream can an ice cream cone hold (within the cone) if its height is 6 inches and the radius of its base is 2.5 inches? Include correct units with your solution. Use "pi" in place of the pi symbol.

OpenStudy (asadkarim7):

cone ' s vol = 1/3 pi r^2 x height = 1/3 x 6.25 x 6 x pi 12.5 pi

OpenStudy (asadkarim7):

12.5 pi in^2

OpenStudy (anonymous):

\[volume = 1/3*\pi*r^2*h\] \[=2*2.5*2.5*\pi\] \[=12.5\pi\]